# Extracting self-similarity from data

### v1

@@
The identification of self-similarity is an indispensable tool for understanding and modelling physical phenomena. Unfortunately, this is not always possible to perform formally in highly complex problems. We propose a methodology to extract the similarity variables of a self-similar physical process directly from data, without prior knowledge of the governing equations or boundary conditions, based on an optimization problem and symbolic regression. We analyze the accuracy and robustness of our method in four problems which have been influential in fluid mechanics research: a laminar boundary layer, Burger’s equation, a turbulent wake, and a collapsing cavity. Our analysis considers datasets acquired via both numerical and wind-tunnel experiments.
@@ Introduction
In this work, we present a generalized methodology that can identify if similarity variables exist and, if they do, their mathematical expressions directly from data, without prior knowledge of the governing equations or boundary conditions. Our work differs from previous efforts (e.g.,  (?, ?)) in that it is not based on dimensional analysis (although it is dimensionally consistent) and can thus identify self-similarities beyond dilational transformations connected to dimensional problems. This is achieved by formulating a minimization problem to identify the similarity variables, which are then interpreted analytically via symbolic regression. In particular, our method simultaneously searches for the optimal non-linear transformations of both independent variables (spatiotemporal coordinates) and dependent variables (observables) that yield the similarity variables of the problem and is thus different to methods aiming to identify coordinate transformations that render hidden symmetries manifest (e.g.,  (?)). In the next section and supplementary material we describe the algorithm that extracts the similarity variables of a self-similar physical process from data. The potential utility of the proposed method is demonstrated by applying it to four problems which have been influential in fluid dynamics research and are exactly or approximately self-similar under different transformation types, with data derived from both laboratory and numerical experiments.
@@ Results
We simulate the collapse of a cavity ($\text{Bo}=10^{-3}$, ${\text{La}}=2500$) by solving the two-phase incompressible axisymmetric Navier-Stokes equations using Basilisk (www.basilisk.fr), which is an open-source flow solver that has been used in computational studies of bursting bubbles  (?, ?, ?). The evolution of the liquid-gas interface is shown in Figs. 4(a-d). We extract profiles of the interface $h(r,t)$ as it approaches the axis of symmetry, at $t_{*}=\left(t_{0}-t\right)/t_{c}=\left[10,9,8,7,6\right]$, where $t_{c}$ is the characteristic time of the horizontal capillary wave  (?), and $t_{0}$ is the moment where the capillary waves meet at the axis of symmetry, $r=0$. The extracted profiles are shown in Fig. 4(e).  (?, ?) observed that in this time window, the interface approximately follows the $\left(t_{0}-t\right)^{2/3}$ scaling of inviscid theory  (?). The algorithm successfully collapses the extracted profiles onto a single curve (Fig. 4(f)) for transformations of the form $\tilde{r}=\alpha(t)r$ for the radial coordinate and $\tilde{h}=\beta(t)h+\gamma(t)$ for the liquid-gas interface. By regressing the identified transformations $\alpha$, $\beta$, and $\gamma$ as power-law functions of the non-dimensional time $t_{*}$ (Fig. 4(g-i)), we retrieve scalings close (apart from an arbitrary multiplicative constant) to the theoretically derived ones  (?, ?, ?, ?, ?).
@@ Discussion
We have presented a methodology for identifying similarity variables from data, in the absence of governing equations. Given measurements of a quantity of interest, the method computes the discrete values of the similarity transformations that best collapse the data and expresses their analytic form via symbolic regression. The self-similarity that we identify is only approximate, in the sense it is the best one that fits the data, and is not derived from the governing partial differential equations and the boundary conditions of the specific problem. The capabilities of the method were demonstrated in four fluid mechanical problems, which accept exact or approximate similarity solutions, based on data from both numerical and laboratory experiments. The transformations identified by the method were in agreement to the analytically or empirically derived similarity transformations, while circumventing the need for scaling analysis or similarity arguments. The present work opens up opportunities for identification of similarity and scaling laws in situations where rigorous mathematical analysis is challenging. This includes a wide range of applications in fluid physics but also in other processes such as quasicrystal shape and growth  (?), stellar collapse  (?), single protein dynamics  (?), and others.
@@ Acknowledgments
Parts of this work were conducted while NB was at Imperial College London, supported by EPSRC, Grant No. EP/W026686/1. LM acknowledges support from EPSRC, Grant Nos. EP/W026686/1 and EP/Y005619/1, the ERC Starting Grant PhyCo No. 949388 and the EU-PNRR YoungResearcher TWIN ERC-PI_0000005.
@@
N.B. and K.S. designed and performed research. N.B., L.M., and K.S. wrote the paper.
@@
The code and input data required to reproduce the work reported in the manuscript are available in the GitHub repository https://github.com/nbeb/extracting_self-similarity_from_data.

### v2

@@
The identification of self-similarity is an indispensable tool for understanding and modelling physical phenomena. Unfortunately, this is not always possible to perform formally in highly complex problems. We propose a methodology to extract the similarity variables of a self-similar physical process directly from data, without prior knowledge of the governing equations or boundary conditions, based on an optimization problem and symbolic regression. We analyze the accuracy and robustness of our method in five problems which have been influential in fluid mechanics research: a laminar boundary layer, Burger’s equation, a turbulent wake, a collapsing cavity, and grid-generated turbulence. Our analysis considers datasets acquired via both numerical and wind-tunnel experiments. The algorithm recovers the known self-similarity expressions in the first four problems and generates new insights in the final problem (grid turbulence).
@@ Introduction
In this work, we present a generalized methodology that can identify if similarity variables exist and, if they do, their mathematical expressions directly from data, without prior knowledge of the governing equations or boundary conditions. Our work differs from previous efforts (e.g.,  (?, ?)) in that it is not based on dimensional analysis (although it is dimensionally consistent) and can thus identify self-similarities beyond dilational transformations connected to dimensional problems. This is achieved by formulating a minimization problem to identify the similarity variables, which are then interpreted analytically via symbolic regression. In particular, our method simultaneously searches for the optimal non-linear transformations of both independent variables (spatiotemporal coordinates) and dependent variables (observables) that yield the similarity variables of the problem and is thus different to methods aiming to identify coordinate transformations that render hidden symmetries manifest (e.g.,  (?)). In the next section and supplementary material we describe the algorithm that extracts the similarity variables of a self-similar physical process from data. The potential utility of the proposed method is demonstrated by applying it to five problems which have been influential in fluid dynamics research and are exactly or approximately self-similar under different transformation types, with data derived from both laboratory and numerical experiments.
@@ Results
We simulate the collapse of a cavity ($\text{Bo}=10^{-3}$, ${\text{La}}=2500$) by solving the two-phase incompressible axisymmetric Navier-Stokes equations using Basilisk (www.basilisk.fr), which is an open-source flow solver that has been used in computational studies of bursting bubbles  (?, ?, ?). The evolution of the liquid-gas interface is shown in Figs. 4(a-d). We extract profiles of the interface $h(r,t)$ as it approaches the axis of symmetry, at $t_{*}=\left(t_{0}-t\right)/t_{c}=\left[10,9,8,7,6\right]$, where $t_{c}$ is the characteristic time of the horizontal capillary wave  (?), and $t_{0}$ is the moment where the capillary waves meet at the axis of symmetry, $r=0$. The extracted profiles are shown in Fig. 4(e).  (?, ?) observed that in this time window, the interface approximately follows the $\left(t_{0}-t\right)^{2/3}$ scaling of inviscid theory  (?). The algorithm successfully collapses the extracted profiles into a single curve (Fig. 4(f)) for transformations of the form $\tilde{r}=\alpha(t)r$ for the radial coordinate and $\tilde{h}=\beta(t)h+\gamma(t)$ for the liquid-gas interface. By regressing the identified transformations $\alpha$, $\beta$, and $\gamma$ as power-law functions of the non-dimensional time $t_{*}$ (Fig. 4(g-i)), we retrieve scalings close (apart from an arbitrary multiplicative constant) to the theoretically derived ones  (?, ?, ?, ?, ?).
Grid turbulence
We consider the case of homogeneous decaying turbulence, experimentally realized by passing a stream of fluid through a uniformly spaced grid inside a wind tunnel (Fig. 5(a)). We analyze measurements of velocity time-series, obtained via hot-wire anemometry at seven positions downstream of the grid (the details of the experimental campaign can be found in Ref.  (?)). The measured turbulence is fully developed, approximately homogeneous and yields the -5/3 law for the energy spectrum at intermediately-sized eddies, as predicted by Kolmogorov’s K41 framework  (?, ?, ?). This configuration puts our algorithm to a real test: Our algorithm can identify self-similarities based on a single scale, while the turbulence cascade is generally accepted to be a two-scale problem.
The theoretical derivation of the -5/3 law assumes that at sufficiently high Reynolds numbers, an intermediate self-similar region forms in the cascade, which is independent of large- and small scale effects  (?, ?, ?). However, the above description is part of a more complex flow-picture, as the widely studied paradigm of grid turbulence sufficiently far from initial conditions shows. The turbulence cascade in that case is generally accepted to be characterized by two self-similarities, both present at the same time and at different eddy sizes: One at large scales where viscosity is negligible  (?, ?), and one at small scales where non-equilibrium effects are negligible  (?, ?). The energy spectrum thus accepts the following general expression  (?)
$E_{11}(\kappa,x)=\epsilon(x)^{2/3}\kappa^{-5/3}f(\kappa L)g(\kappa\eta)\,,$ (9)
where $x$, $\kappa$, $E_{11}(\kappa,x)$, $\epsilon(x)$ represent the distance from the grid, wavenumber, energy spectrum and dissipation rate, respectively. $L$ and $\eta$ are the integral and Kolmogorov scales, characteristic of the large- and small-scale self-similarities, respectively. To retrieve the -5/3 law, $E_{11}(k,x)=\epsilon(x)^{2/3}\kappa^{-5/3}$, one needs to be asymptotically far from large scales (i.e., $\kappa L\to\infty$) and far from small scales ($\kappa\eta\to 0$) at the same time, as in that case K41 assures that both $f$ and $g$ tend to unity. It is of interest to see how our single-scale algorithm fares in this two-scale problem.
Figure 5(b) plots the measured power spectral densities versus the wavenumbers at different measurement stations. Figures 5(c,d) show the collapse of the power spectral densities using large and small scale similarity variables, respectively. Using a library that consists of $(L,\eta)$ for the $x$ axis and $(\epsilon,L,\eta,k)$ for the $y$ axis, symbolic regression of the algorithmically identified stretching transformations yields the following expressions, plotted in Fig. 5(e)
$\tilde{\kappa}\propto\kappa L^{0.368}\eta^{0.632},\qquad\tilde{E_{11}}\propto E_{11}\epsilon^{0.695}L^{0.709}\eta^{-1.015}k^{-2.042}.$ (10)
As expected, our algorithm fails to derive the two-scale similarity expression (9). Further insight may be obtained by using the dissipation scaling $\epsilon\propto k^{3/2}/L$, where $k$ is the turbulence kinetic energy, which is known to characterize homogeneous decaying turbulence far from initial conditions  (?, ?, ?). Expression (10) then becomes
$\tilde{\kappa}\propto\kappa L^{0.368}\eta^{0.632},\qquad\tilde{E_{11}}\propto E_{11}\epsilon^{-0.667}\left(L^{0.391}\eta^{0.609}\right)^{-5/3}.$ (11)
Inspection of the above expression reveals that our algorithm has produced novel information by extracting self-similarity based on a single empirical length scale, $l\approx L^{0.38}\eta^{0.62}$, which is a combination of $L$ and $\eta$. It is very interesting that a single length scale theory for the cascade does in fact exist (see Ref.  (?)). The single length scale of that theory is the so-called Taylor microscale, which in our problem becomes $\lambda=L^{1/3}\eta^{2/3}$  (?), i.e., very close to our empirically-derived length scale. Therefore, while our algorithm may not be appropriate for the extraction of two-scale self-similarities, it is in fact able to contribute to the discussion regarding the modeling and behaviour of the turbulence cascade.
Figure 5: Data-driven identification of self-similarity in decaying turbulence. (a) Schematic of the experimental set-up. (b) Experimentally measured power spectral densities. The dashed vertical line delineates the range of the spectrum that is used. (c) Measured spectrum normalised by the inertial scales. (d) Measured spectrum normalised by the Kolmogorov scales. (e) Measured spectrum normalised by the algorithmically identified expression (Eq. (10)).
@@ Discussion
We have presented a methodology for identifying similarity variables from data, in the absence of governing equations. Given measurements of a quantity of interest, the method computes the discrete values of the similarity transformations that best collapse the data and expresses their analytic form via symbolic regression. The self-similarity that we identify is only approximate, in the sense that it is the best one that fits the data, and is not derived from the governing partial differential equations and the boundary conditions of the specific problem. The capabilities of the method were demonstrated in five fluid mechanical problems, which are known or assumed to accept similarity solutions, based on data from both numerical and laboratory experiments. In four problems, the transformations identified by the method were in agreement to the analytically or empirically derived similarity transformations, while circumventing the need for scaling analysis or similarity arguments. In the fifth problem (turbulence cascade) our method derived a novel expression for the energy spectrum, different from the widely accepted two-scale similarity expressions, contributing in that way to the discussion regarding the nature of similarity in turbulence and showcasing how the present work opens up opportunities for identification of similarity and scaling laws in situations where rigorous mathematical analysis is challenging. This includes a wide range of applications in fluid physics but also in other processes such as quasicrystal shape and growth  (?), stellar collapse  (?), single protein dynamics  (?), and others.
@@ Acknowledgments
Parts of this work were conducted while NB was at Imperial College London, supported by EPSRC, Grant No. EP/W026686/1. LM acknowledges support from EPSRC, Grant Nos. EP/W026686/1 and EP/Y005619/1, the ERC Starting Grant PhyCo No. 949388 and the EU-PNRR YoungResearcher TWIN ERC-PI_0000005. KS acknowledges support from the ERC Starting Grant ONSET No. 101163321.
@@
N.B., L.M. and K.S. designed research; N.B. and K.S. performed research; N.B. analyzed data; and N.B. and K.S. wrote the paper.
@@
The code and input data required to reproduce the work reported in the manuscript are available in the GitHub repository https://github.com/nbeb/extracting_self-similarity_from_data. The grid turbulence data are available upon request.